how do I solve |x-9|=|x+6|?
this is what I did:
x-9=x+6
x-9=6
x=15
x-9=x+6
x-9=-x-6
2x=3
x=1
1,15 (not thinking this is correct...)

Answer by checkley77(12844)

|x-9|=|x+6|
x-9=-(x+6)
x-9=-x-6
2x=-6+9
2x=3
x=3/2 ans.
-(x-9)=x+6
-x+9=x+6
-x-x=6-9
-2x=-3
x=-3/-2
x=3/2 ans.